OpenStudy (anonymous):

What are the x– and y–intercepts of the quadratic equation y = (x – 3)^2– 25? Identify the x–intercepts by graphing the quadratic equation.

OpenStudy (mimi_x3):

y - intercept , x=0 y = - 16 x-intercept , y = 0 (x-3)^2 - 25 = 0 x^2 - 6x + 9 - 25 = 0 x^2 -6x - 16 = 0 (x-8)(x+2) = x = 8, -2